REU 2007: SCHEDULE FOR WEEKS ONE AND TWO

Morning, TTh 1:30, and M-F 4:00 talks: Ryerson 251 1:30 MWF talks: Eckhardt 133

Talks may not take up the full amount of time scheduled: maxima are listed

MONDAY, June 18:

Welcome and opening meeting: 9:00 in RYERSON 251. Mentoring organizational meeting: 11:30 in RYERSON 251. Pizza and coke: 12:30 in the barn.

FRIDAY, June 22:

1:00 meeting in the barn for those working in YSP and SESAME. FIRST WEEK:

Date Babai Farb Lawler May
Mon, June 18 9:15–11:15  1:30-3:00 4:00-5:30
Tue, June 19 9:00–12:00 1:30–3:30
Wed, June 20 9:00–12:00  1:30-3:00 4:00-5:30
Thu, June 21 9:00–12:00 1:30–3:30
Fri, June 22 9:00–12:00  1:30-3:00 4:00-5:30

TUTORIALS: 3:30-6:00 TTh, lunchtimes, evenings SECOND WEEK: NOTE: attendance at the YSP and SESAME morning preparation and training sessions by all people who will be counsellors in either of these programs is required.

Date  YSP/SES  Babai  Babai  Farb  Lawler  May
Apprentice Discrete
Mon, June 25 9:00–12:00 9:00–12:00  1:30–3:00 4:00–5:30
Tue, June 26 9:00–12:00 9:00–12:00 4:00–5:30 1:30–3:30
Wed, June 27 9:00–12:00 9:00–12:00  1:30–3:00 4:00–5:30
Thu, June 28 9:00–12:00 9:00–12:00 4:00–5:30 1:30–3:30
Fri, June 29  9:00–12:00 9:00–12:00  1:30–3:00 4:00–5:30

TUTORIALS: lunchtimes, evenings

Here is the schedule in summary form:

FIRST WEEK:

Babai (Discrete mathematics): 9:15–11:15 M, 9:00–12:00 T-Fr Farb (Mapping class groups): 1:30–3:30 TTh Lawler (Random walks; Brownian motion) 1:30–3:00 MWF May (Surfaces, categories, algebras, TQFT’s): 4:00–5:30 MWF

SECOND WEEK:

YSP/SESAME Preparation: 9:00–12:00 M–F Babai (Apprentice): 9:00–12:00 M–F Babai (Discrete mathematics): 4:00–5:30 TTh Farb (Mapping class groups): 1:30–3:30 TTh Lawler (Random walks; Brownian motion) 1:30–3:00 MWF May (Surfaces, categories, algebras, TQFT’s): 4:00–5:30 MWF